OpenStudy (anonymous):

PLEASE HELP I WILL GIVE MEDAL! Which of these ordered pairs is a solution to the inequality 2x + y > – 4? (–1, –1) (0, –5) (4, –12) (–3, 0)

OpenStudy (mathstudent55):

Replace x and y by each pair of values and calculate.

OpenStudy (anonymous):

plug each of those values in and see which one works so: 2 x (-1) + (-1) 2 x (0) + (5) 2 x (4) + (-12) 2 x (-3) + (0) Solve those equations

OpenStudy (anonymous):

1. -2x-1

OpenStudy (anonymous):

2. 5

OpenStudy (anonymous):

oops, b. is supposed to be 2 x (0) + (-5)

OpenStudy (anonymous):

so 2. is -5

OpenStudy (anonymous):

yes ?? i think

OpenStudy (anonymous):

here are the answers to the equations: a. 2 x (-1) + (-1) = -3

OpenStudy (anonymous):

b. 2 x (0) + (-5) = -5

OpenStudy (anonymous):

3. 8x-12

OpenStudy (anonymous):

c. 2 x (4) + (-12) = -4

OpenStudy (anonymous):

4. -6

OpenStudy (anonymous):

d. 2 x (-3) + (0) = 6

OpenStudy (anonymous):

right, so we have: a. -3 b. -5 c. -4 d. -6

OpenStudy (anonymous):

wooow . way off with my answers

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so which one fits the equation > -4? Which number is greater than -4?

OpenStudy (anonymous):

oh i got 1 correct out of the 4 you gave me !! i suck at math . like seriously

OpenStudy (anonymous):

Lol, it's okay, just make sure you think through what you need to get out of your equation before you start it.

OpenStudy (anonymous):

lol ok . sorry . but a number greater than -4 can be a positive right ..

OpenStudy (anonymous):

it can be positive, but remember you only have negative numbers, set up a number line: -6 -5 -4 -3 -2 -1 0 ________l________l________l_______l________l________l________l

OpenStudy (anonymous):

your answers are: a. -3 b. -5 c. -4 d. -6

OpenStudy (anonymous):

So which one is greater than -4?